OpenStudy (phoenixfire):

Use elementary row operations to find the inverse of \left[\begin{matrix}-2 & -1 & 3 \\ 4 & -2 & 1 \\ 0 & 4 & -2\end{matrix}\right] I know the inverse as I've done it via cofactor and adjugate. And wolframalpha has confirmed the result. But no matter how many times I try with the elementary row operations I can't get it right.

OpenStudy (unklerhaukus):

\[\qquad\qquad\quad\quad[\textbf A|\textbf I]\]\[\left[\begin{matrix}-2 & -1 & 3 \\ 4 & -2 & 1 \\ 0 & 4 & -2\end{matrix}\right|\left.\begin{matrix}1 & 0 & 0\\ 0 & 1 & 0 \\ 0 &0 & 1\end{matrix}\right]\]

OpenStudy (unklerhaukus):

\[\left[\begin{matrix}-2 & -1 & 3 \\ 0 & -4 & 7 \\ 0 & 4 & -2\end{matrix}\right|\left.\begin{matrix}1 & 0 & 0\\ 2 & 1 & 0 \\ 0 &0 & 1\end{matrix}\right]\qquad R_2\rightarrow R_2+2R_1\]

OpenStudy (unklerhaukus):

\[\left[\begin{matrix}-2 & -1 & 3 \\ 0 & -4 & 7 \\ 0 & 0& 5 \end{matrix}\right|\left.\begin{matrix}1 & 0 & 0\\ 2 & 1 & 0 \\ 2 &1 & 1 \end{matrix}\right]\qquad R_3\rightarrow R_3+R_2\]

OpenStudy (unklerhaukus):

\[\left[\begin{matrix}2 & 1 & 2 \\ 0 & -4 & 7 \\ 0 & 0& 5 \end{matrix}\right|\left.\begin{matrix}1 & 1 & 1\\ 2 & 1 & 0 \\ 2 &1 & 1 \end{matrix}\right]\qquad R_1\rightarrow -(R_1-R_3)\]

OpenStudy (unklerhaukus):

\[\left[\begin{matrix}2 & 1 & 2 \\ 0 & 4 & -2 \\ 0 & 0& 5 \end{matrix}\right|\left.\begin{matrix}1 & 1 & 1\\ 0 & 0 & 1 \\ 2 &1 & 1 \end{matrix}\right]\qquad R_2\rightarrow -(R_2-R_3)\]

OpenStudy (unklerhaukus):

\[\left[\begin{matrix}2 & 1 & 2 \\ 0 & 1 & -1/2 \\ 0 & 0& 5 \end{matrix}\right|\left.\begin{matrix}1 & 1 & 1\\ 0 & 0 & 1/4 \\ 2 &1 & 1 \end{matrix}\right]\qquad R_2\rightarrow R_2/4\]

OpenStudy (unklerhaukus):

\[\left[\begin{matrix}2 & 0& 5/2 \\ 0 & 1 & -1/2 \\ 0 & 0& 5 \end{matrix}\right|\left.\begin{matrix}1 & 1 & 3/4\\ 0 & 0 & 1/4 \\ 2 &1 & 1 \end{matrix}\right]\qquad R_1\rightarrow R_1-R_2\]

OpenStudy (unklerhaukus):

\[\left[\begin{matrix}1 & 0& 5/4 \\ 0 & 1 & -1/2 \\ 0 & 0& 5 \end{matrix}\right|\left.\begin{matrix}1/2 & 1/2 & 3/8\\ 0 & 0 & 1/4 \\ 2 &1 & 1 \end{matrix}\right]\qquad R_1\rightarrow R_1/2\]

OpenStudy (unklerhaukus):

\[\left[\begin{matrix}1 & 0& 5/4 \\ 0 & 1 & -1/2 \\ 0 & 0& 1 \end{matrix}\right|\left.\begin{matrix}1/2 & 1/2 & 3/8\\ 0 & 0 & 1/4 \\ 2/5 &1/5 & 1/5 \end{matrix}\right]\qquad R_3\rightarrow R_3/5\]

OpenStudy (unklerhaukus):

\[\left[\begin{matrix}1 & 0& 5/4 \\ 0 & 1 & 0 \\ 0 & 0& 1 \end{matrix}\right|\left.\begin{matrix}1/2 & 1/2 & 3/8\\ 1/5 & 1/10 & 7/20 \\ 2/5 &1/5 & 1/5 \end{matrix}\right]\qquad R_2\rightarrow R_2+R_3/2\]

OpenStudy (unklerhaukus):

\[\left[\begin{matrix}1 & 0& 0 \\ 0 & 1 & 0 \\ 0 & 0& 1 \end{matrix}\right|\left.\begin{matrix}0 & 1/4 & 1 /8\\ 1/5 & 1/10 & 7/20 \\ 2/5 &1/5 & 1/5 \end{matrix}\right]\qquad R_1\rightarrow R_1-5 R_3/4\]\[\qquad\qquad[\textbf I |\textbf A^{-1}]\]

OpenStudy (anonymous):

wow that is a lot of work^

OpenStudy (unklerhaukus):

\[\color{gray} {\begin{matrix}0 & & 0 \\ & & \\ & \smile & \end{matrix}} \]

OpenStudy (phoenixfire):

Wow thank you @UnkleRhaukus !! I understand why you make below the diagonal 0, but then you started working on the top triangle. Does the order matter? For example can I make the bottom triangle 0 then work down the diagonal making them 1? Is making the bottom triangle 0 the really only required step to be done first?

OpenStudy (unklerhaukus):

you can perform any of the allowed row operations at any time and in any order, some ways are certainly quicker than others , im not sure if the steps i used were the simplest, but they did work

OpenStudy (phoenixfire):

Okay, thank you again.

OpenStudy (unklerhaukus):

http://www.math.odu.edu/~bogacki/cgi-bin/lat.cgi